OpenStudy (anonymous):

Parametric equation with calculus help. I am supposed to calculate the speed of an object that has these two parametric equations x(t)= e^t + e^-t y(t) = e^-t

zepdrix (zepdrix):

\[\Large\rm \frac{dy}{dx}=\frac{(dy/dt)}{(dx/dt)}\]

zepdrix (zepdrix):

And then speed is uhhhh, directionless, yes? So we need the magnitude of whatever velocity we get.

zepdrix (zepdrix):

So looks like we need to take a couple derivatives :o

OpenStudy (anonymous):

I tried that and I got derivative that ended up looking like this: dy/dt = (-e^-t)/(e^t -e^-t). However I'm not getting the right answers even though I did the exact same thing you brought up.

OpenStudy (anonymous):

Damnit I mean dy/dx not dy/dt

zepdrix (zepdrix):

\[\Large\rm \frac{dy}{dx}=\frac{(dy/dt)}{(dx/dt)}=\frac{-e^{-t}}{e^t-e^{-t}}\]And that's not working out? Let's um.. try to clean it up a bit maybe. I dunno.

OpenStudy (anonymous):

yeah that's exactly what I got

OpenStudy (anonymous):

Doesn't work though so idk.

zepdrix (zepdrix):

\[\Large\rm \frac{-e^{-t}}{e^t-e^{-t}}\color{royalblue}{\left(\frac{e^t+e^{-t}}{e^t+e^{-t}}\right)}=\frac{e^{-2t}-1}{e^{2t}-e^{-2t}}\]Ehh I guess that doesn't look at that much better lol. Maybe we just need to take the negative off the front, to show that a SPEED should be directionless, no negative direction.\[\Large\rm \frac{e^{-t}}{e^t-e^{-t}}\]Ahhh I dunno >.< Hmmm

OpenStudy (anonymous):

Hmm that's actually not a bad idea because in physics speed is a scalar not a vector, meaning it can't be negative.

OpenStudy (anonymous):

The question is asking for speed, but it's for a calculus class so idk maybe they're assuming the student knows that speed is a scalar?

zepdrix (zepdrix):

I know that when you get into multivariable calculus you're expected to know that speed is a scalar. I can't remember before that though.

zepdrix (zepdrix):

You're trying to enter this into a site or something?

OpenStudy (anonymous):

Yeah exactly.

OpenStudy (nincompoop):

yeah in physics 1, speed is scalar, and velocity is vector

OpenStudy (anonymous):

yeah exactly nincompoop.

OpenStudy (nincompoop):

in math, they don't give a flip

OpenStudy (nincompoop):

you define them

OpenStudy (anonymous):

It's the only thing I can think of that would make sense, I mean I did the (dy/dt)/(dx/dt) thng and I got the exact same equation zepdrix did.

zepdrix (zepdrix):

Have you entered other problems without having any issues? Like you're comfortable with the input that you need for whatever site you're using? Sometimes they expect exp{} for the exponential base e.

OpenStudy (anonymous):

For the most part yes, but sometimes the website can be dumb so it could just be the website. However there is another problem that I need to raise with the equation.

OpenStudy (anonymous):

The problem I am tackling is specifically about calculating minimums and maximums on the interval [0,3] but as you can see with the equation we derived, if you set t = 0 the derivative does not exist at that point because the denominator is 0. I even calculated the critical numbers for the derivative and got 0 as the only critical number. So that's another problem.

OpenStudy (nincompoop):

what website?

OpenStudy (anonymous):

erm, idk I found this. not sure if it's the right equation?

OpenStudy (anonymous):

webassign.

zepdrix (zepdrix):

I did some google stuff... are we supposed to use this instead for parametric?\[\Large\rm v(t)=\sqrt{\left(\frac{dx}{dt}\right)^2+\left(\frac{dy}{dt}\right)^2}\]

OpenStudy (nincompoop):

i knew it laughing out loud

OpenStudy (anonymous):

except replace v(t) with dp(t) :))

OpenStudy (anonymous):

Idk possibly.

OpenStudy (anonymous):

We never learned that though so that would be odd if that were the case.

OpenStudy (nincompoop):

arc length uses pythagorean

OpenStudy (anonymous):

?

OpenStudy (anonymous):

I'll try this new equation and see if it works.

OpenStudy (nincompoop):

look at the three key points, and you will be set for life

OpenStudy (anonymous):

Lol that equation worked, the Pythagorean one.

OpenStudy (anonymous):

Thanks guys, strange though because we were never taught this but I'm going to save the pages you gave me and study them.

OpenStudy (anonymous):

Idk who I shall give my medal too.....

OpenStudy (anonymous):

I'll flip a coin heads goes to zepdrix, tails goes to nincompoop

OpenStudy (anonymous):

Zepdrix wins

OpenStudy (nincompoop):

I don't want medals a masters degree is enough for me to aim for

OpenStudy (anonymous):

Ha, thanks for the help I'm going to close this now.